# GLL: A Differentiable Graph Learning Layer for Neural Networks

### v1

@@ 4.2 Comparison to the Multilayer Perceptron (MLP)
    Smooth training and improved convergence with GLL: In Figure 4, GLL-0, trained from scratch, achieves good accuracy and low loss function in a short time, and the performance continues to improve slowly as the training progresses (up to 50 epochs). It is worth noting that the training of MLP quickly stagnates, with the loss and accuracy showing little change over a long period. However, the training continues very smoothly when we replace the MLP training process with GLL midway, i.e., GLL-50 and GLL-75. One possible explanation of this phenomenon is if the number of parameters or the number of layers is large, it may lead to the dying ReLU problem (Lu et al. (2020)) or vanishing gradients (Hochreiter (1998)), causing training stagnation. On the other hand, since GLL does not contain any trainable parameters, it can make the optimization process more smooth.
@@ 4.3 Adversarial Robustness
In this subsection, we demonstrate our method’s superior robustness to adversarial attacks compared to the typical projection head and softmax activation layer in a neural network. Adversarial attacks seek to generate adversarial examples (Szegedy et al. (2013)): input data with small perturbations that cause the model to misclassify the data. Many highly performant deep neural networks can be fooled by very small changes to the input, even when the change is imperceptible to a human (Madry et al. (2017); Goodfellow et al. (2014)). Hence, training models robust (i.e. resistant) to such attacks is important for both security and performance. Our experiments show that in both naturally and robustly-trained models, using the graph learning layer improves adversarial robustness compared to a softmax classifier on MNIST (Yann LeCun (2010)), FashionMNIST (Xiao et al. (2017)), and CIFAR-10 (Krizhevsky et al. (2009)) in both relatively shallow and deep networks. Moreover, GLL is more robust to adversarial attacks without sacrificing performance on natural images.
@@ 4.3.1 Adversarial Attacks
We consider three types of adversarial attacks: the fast gradient sign method (FGSM) (Goodfellow et al. (2014)) in the $\ell_{\infty}$ norm, the iterative fast gradient sign method (IFGSM) (Kurakin et al. (2018)) in the $\ell_{\infty}$ norm, and the Carlini-Wagner attack (Carlini and Wagner (2017)) in the $\ell_{2}$ norm.
@@
where $x_{0}=x$ and $i=0,...,N-1$, and so $x_{N}$ is the generated adversarial example. However, we must also clip the iterates (pixel-wise) during this iteration to ensure that (1) they remain in the $\epsilon$-neighborhood of $x$ and (2) they remain in the range of the image space (e.g. in [0,1] for a grayscale image). Thus, the IFGSM attack is:
@@
Note that this assumes the pixel values are in the range $[0,1]$, but can be easily tailored to other settings. IFGSM attacks have been shown to be more effective than FGSM attacks (Kurakin et al. (2018)), however we will see in Section 4.3.3 that our graph learning layer is highly robust to this iterative method.
@@
The constraint $C(x+\delta)=t$ is highly non-linear, so Carlini and Wagner instead consider objective functions $g$ that satisfy $C(x+\delta)=t$ if and only if $g(x+\delta)\leq 0$. While there are many choices for $g$ (see Carlini and Wagner (2017), Section IV A), we will use the following in our experiments:
@@
where $a^{+}=\max(0,a)$. It is worth noting that a common choice for $g$ in the literature is to use the unnormalized logits from the neural network (that is, the output of a neural network just before the standrad softmax layer) instead of $F(x)$. While both definitions satisfy the condition, we use the above because graph learning methods (such as Laplace learning) operate in the feature space and directly output probability distributions over the classes, so there is no notion of logits.
@@
where $c>0$ is a suitably chosen hyperparameter. To make the problem unconstrained, we can introduce a variable $w$ and write $\delta=\frac{1}{2}(\tanh{w}+1)-x$; since $-1\leq\tanh{w}\leq 1$, we have that $0\leq x+\delta\leq 1$. This gives the unconstrained objective:
@@
which can be solved using standard optimization routines; we use the Adam optimizer (Kingma and Ba (2014)).
@@ 4.3.2 Adversarial Training
To train adversarially robust networks, we use projected gradient descent (PGD) adversarial training (Madry et al. (2017)) to our models. Motivated by guaranteeing robustness to first-order adversaries, PGD training applies IFGSM to images with an random initial perturbation during training. Hence, the model is trained on adversarial examples. More precisely, PGD training consists of two steps each batch: first, $x$ is randomly perturbed
@@
where $x_{0}=x^{*}$ and the number of iterations and $\alpha$ are hyperparameters. This training procedure, while more expensive, has been shown to be effective at improving robustness in deep learning models compared to training on clean (i.e. natural) images (Madry et al. (2017), Wang and Osher (2021)).
@@ 4.3.3 Adversarial Experiments
We evaluate the adversarial robustness of neural networks with the GLL classifier on the MNIST (Yann LeCun (2010)), FashionMNIST (Xiao et al. (2017)), and CIFAR-10 (Krizhevsky et al. (2009)) benchmarks datasets, and compare our method to the same networks with a softmax classifier. For all datasets, we normalize the images to have mean 0 and standard deviation 1, and we train and run all adversarial attacks on a 11GB GeForce RTX 2080Ti GPU. We note that the goal of these experiments is not to get state-of-the-art natural accuracies; while our natural accuracy on these benchmark datasets is strong, the purpose of these experiments is to demonstrate our method’s superior robustness to adversarial attacks compared to the typical classifier (a linear layer and softmax layer), on a variety of networks and datasets, without sacrificing performance on natural images. A summary of our experiments - highlighting a handful of our adversarial robustness results - can be found in Table 4, while complete results are shown in Figures 5, 6, and 7. Throughout, when we say “MLP” we mean a linear layer mapping from the feature space to the logits followed by a softmax classifier, while “GLL” refers to our graph learning layer classifier.
@@
MNIST GL Robust 99.45 97.72 97.70 98.84
MLP Robust 99.20 95.17 93.57 97.41
GL Natural 99.53 96.51 96.44 98.71
MLP Natural 99.25 94.26 91.52 97.24
Fashion-MNIST GL Robust 91.04 88.07 87.17 83.81
MLP Robust 92.08 86.50 85.67 81.00
GL Natural 92.14 65.44 61.81 57.19
MLP Natural 91.65 40.81 18.30 62.01
CIFAR-10 GL Robust 90.60 80.50 79.62 78.09
MLP Robust 91.29 73.95 71.04 75.94
GL Natural 94.29 59.05 54.41 77.23
MLP Natural 94.98 50.75 8.21 53.00
@@
Table 4: Summary of adversarial results of our method compared to the standard linear layer-Softmax classifier, on both robustly and naturally trained models (column 2). We report natural accuracy (i.e. no attacks), and adversarial accuracy on FGSM, IFGSM, and CW attacks. For MNIST, we let $\epsilon=0.3$ for the FGSM and IFGSM attacks, and $\epsilon=0.05$ for the other datasets. We report CW results for $c=20$. We report more complete results across a range of values for $\epsilon$ and $c$ in Figures 5, 6, and 7. We include the natural accuracies to show that the GLL layer does not sacrifice performance in this regime, while also being significantly more robust to adversaries. The slight drop in natural accuracy on robustly trained models compared to naturally trained models has been observed previously in Madry et al. (2017); however, as expected, the robust models perform significantly better on adversarially-perturbed data than their naturally-trained counterparts.
@@
For MNIST, we use the same “small CNN” architecture as in Wang and Osher (2021): four convolutional layers with ReLU activations after each layer and maxpooling after the 2nd and 4th layer, followed by 3 linear layers and a softmax classifier. To adopt this CNN to our approach, we replace the final linear layer and softmax with our GLL. We train for 100 epochs using the Adam optimizer (Kingma and Ba (2014)) with an initial learning rate of $0.01$ which decays by a factor of $0.1$ every 25 epochs. For the robustly trained models, we run PGD training with $\epsilon=0.3$ and $\alpha=0.01$ for 5 iterations. In training, we use a batch size of 1000, and for GLL we use 100 labeled “base” points per batch (i.e. 10 per class), which are randomly sampled each epoch.
@@
To evaluate experimental robustness, we attack the test set with the FGSM, IFGSM, and CW attacks. When using FGSM and IFGSM on GLL, we embed the entire test set of 10000 digits along with 10000 randomly selected data points from the training set (i.e. 1000 per class) when running the attacks, while we run CW in batches of 1000 test and 1000 training points (this was simply due to computational constraints; in general graph learning methods perform better with more nodes on the graph as it better approximates the underlying data manifold). For the models with MLP classifier, we use the same batch sizes (full batch on FGSM and IFGSM, 1000 for CW). We report results for FGSM and IFGSM at varying levels of $\epsilon$, fixing $\alpha=0.05$ for IFGSM and letting the number of iterations be $5*(\epsilon/\alpha)$, following a similar heuristic as in Kurakin et al. (2018) to allow the adversarial example sufficient ability to reach the boundary of the $\epsilon$-ball. For CW attacks, we let $c$ vary and run 100 iterations of the Adam optimizer with a learning rate of 0.005. Results for each attack, comparing both robustly and naturally trained models with MLP and GLL classifiers, are shown in Figure 5.
For FashionMNIST, we use ResNet-18 (He et al. (2016a)), where again we replace the final linear layer and softmax with GLL. We again use the Adam optimizer for 100 epochs, with an intial learning rate of $0.01$ which decays by a factor of $0.5$ every 10 epochs. For the robustly trained models, we run PGD training with $\epsilon=0.05$ and $\alpha=0.01$ for 5 iterations. In training, we use a batch size of 2000, and for GLL we use 200 labeled “base” points per batch. We attack in batches of 250 test points, with 500 training points for GLL; this was chosen as it was the largest batch size we could use given our computational resources. The hyperparameters for the attacks are identical to those for MNIST. We report results in Figure 6.
Finally, for CIFAR-10, we use PreActResNet-18 (He et al. (2016b)), where again we replace the final linear layer and softmax with GLL. We use stochastic gradient descent with momentum set to $0.9$, a weight decay of $0.0005$, and an intial learning rate of $0.1$. We train for 150 epochs and use cosine annealing to adjust the learning rate (Loshchilov and Hutter (2016)). For the robustly trained models, we run PGD training with $\epsilon=0.05$ and $\alpha=0.01$ for 5 iterations. In training, we use a batch size of 200, and for GLL we use 100 labeled “base” points per batch. We attack in batches of 200 test points, with 500 training points for GLL; this was chosen as it was the largest batch size we could use given our computational resources. The hyperparameters for the attacks are identical to those for MNIST, except we use 50 iterations of optimization for the CW attacks. We report results in Figure 7.
@@
Figure 5: Adversarial robustness to FGSM, IFGSM, and CW attacks on the MNIST dataset. The backbone network is the same in all cases; the only difference is the classifier (either MLP with softmax or our novel GLL). We report results for both robustly and naturally trained models.
@@
Figure 6: Adversarial robustness to FGSM, IFGSM, and CW attacks on the FashionMNIST dataset. The backbone network is the same in all cases (ResNet-18); the only difference is the classifier (either MLP with softmax or our novel GLL). We report results for both robustly and naturally trained models.

### v2

@@ 4.3 Toy Example: Over-parameterized CNN on FashionMNIST
    Smooth training dynamics. When training starts with an MLP classifier and then switches to GLL (GLL-50 and GLL-75), the loss immediately resumes decreasing and the test accuracy improves further. This shows that GLL can effectively unlock stalled training and provide a smoother optimization trajectory for the shared encoder.
4.4 Large Scale Experiments
Network Params
Encoder MLP head Total
VGG11 9,223,232 332,778 9,556,010
VGG13 9,407,936 332,778 9,740,714
ResNet20 271,824 16,938 288,762
ResNet32 466,256 16,938 483,194
ResNet44 660,688 16,938 677,626
ResNet56 855,120 16,938 872,058
ResNet110 1,730,064 16,938 1,747,002
ResNet18 11,168,832 332,778 11,501,610
PreActResNet18 11,171,146 332,778 11,503,924
Table 3: Parameter counts by network
Train Num Whole (50000) 20% (10000) 2% (1000)
Network MLP WNLL GLL MLP WNLL GLL MLP WNLL GLL
[…]
Table 4: CIFAR-10 test error rates (%, lower is better) comparing MLP, WNLL, and GLL output heads across architectures, evaluated under three training-set sizes: “Whole” (50000 images), “20%” (10000), and “2%” (1000). Entries report mean $\pm$ one standard deviation of the test error. Best results for each backbone and label rate are bolded and results within a standard deviation of the best result are underlined. Network parameter counts follow Table 3.
Train Num Whole (112800) 20% (22560) 2% (2256)
Network MLP WNLL GLL MLP WNLL GLL MLP WNLL GLL
[…]
Table 5: EMNIST test error rates (%, lower is better) comparing MLP, WNLL, and GLL output heads across architectures, evaluated under three training-set sizes: “Whole” (112800 images), “20%” (22560), and “2%” (2256). “Whole” uses the full standard training set. Best results for each backbone and label rate are bolded and results within a standard deviation of the best result are underlined. Entries report mean $\pm$ one standard deviation of the test error. Network parameter counts follow Table 3.
Train Num Whole (112800) 20% (22560) 2% (2256)
Network MLP WNLL GLL MLP WNLL GLL MLP WNLL GLL
[…]
Table 6: EMNIST test error rates (%, lower is better) for MLP, WNLL, and GLL output heads across architectures, evaluated under three training-set sizes: “Whole” (112800 images), “20%” (22560), and “2%” (2256). “Whole” uses the full standard training set. Error rates are computed on a class subindex of the EMNIST (balanced) label space: although the dataset contains 47 classes in total, we evaluate only the 10 hardest classes: 0, 1, 9, F, I, L, O, f, g, and q. Entries report mean $\pm$ one standard deviation of the test error. Best results for each backbone and label rate are bolded and results within a standard deviation of the best result are underlined. Network parameter counts follow Table 3.
This section presents large scale experiments that compare three interchangeable classifiers under the same backbone feature encoder: (1) a two-layer MLP with softmax, (2) WNLL (wang2021graph), and (3) the proposed GLL. We chose (1) as a comparison point because GLL is intended to be a drop-in replacement for the standard MLP-based classification head, while (2) serves as a good intermediary comparison between MLP and GLL because it uses a graph-based classifier in testing. However, WNLL differs significantly from GLL in training. First, it uses an alternating two-stage training protocol, where Stage 1 consists of normal MLP training, and Stage 2 is WNLL training. During Stage 2, the graph classification head computes a loss, but backpropagation is handled by approximating the gradients with a linear layer. Second, Stage 1 lasts for 400 epochs at a time, while Stage 2 lasts for 5 epochs at a time. Finally, Stage 2 freezes the parameters of all but two layers of the network. In this sense, WNLL is closer to MLP in training than GLL.
As discussed in Section 4.1, all models are first warmed up with contrastive pretraining on the backbone without any classification head. Then, we proceed with supervised end-to-end training after attaching one of the three heads. We report the test error on CIFAR-10 and EMNIST-Balanced (cohen2017emnist) datasets. EMNIST extends MNIST by adding handwritten upper and lowercase letters. The Balanced version contains 131,600 samples and 47 classes [^5] Some of the upper and lowercase letters are merged into one class because they are the same shape, for example “s” and “S”.. Following a similar experimental protocol as wang2021graph, we evaluate multiple feature encoders including variants of VGG (simonyan2014very), ResNet (he2016deep), and PreActResNet (he2016identity) and three label fractions: full training set, 20 percent, and 2 percent based on class-balanced sampling, since graph-based classifiers are a standard choice in semi-supervised settings. Parameter counts for these networks are listed in Table 3.
For CIFAR-10, we train for 1000 epochs. Recalling the definitions from Section 4.1, we use a batch of $1500$ points in training and testing, which includes 250 base labeled points $\mathcal{L}_{b}$ for GLL. The remaining 1250 points are constructed with labeled ($\mathcal{B}_{l}$) and unlabeled ($\mathcal{B}_{u}$) points according to Equation 52. This batch partition also applies to WNLL during WNLL’s Stage 2 ($\mathcal{L}_{b}$ is referred to as “template” labels in wang2021graph). We use SGD with a learning rate of $0.001$, momentum of $0.9$, and weight decay of $0.0005$, with cross entropy loss. We use the standard implementations of all the neural network backbones. On EMNIST, we use similar training settings, except we train for 200 epochs and use a learning rate of $0.01$, and the base sample size is increased to 470 to account for the additional number of classes. All experiments were run on a single A100 GPU. For both datasets, we use the data augmentation strategy described in Section 4.1.
Test error rates for CIFAR-10 and EMNIST across architectures and supervision levels are given in Tables 4 and 5. On CIFAR-10, the GLL-based models are consistently the best performing of the three classification heads, especially as training set size decreases. This empirically validates the advantage of the relational information encoded and exploited via a GLL, as well as the precise tracking of gradients. This becomes even more essential for lower label rates.
On EMNIST, GLL is better than WNLL across label rates and MLP for low label rates, and on par with MLP in the other training settings. As expected, the GLL-based networks see the largest advantage in the semi-supervsied setting (2%).
During our experiments on EMNIST, we found that the majority of the classes were easily classified by all methods, while a handful of other classes were consistently more challenging. By isolating results on only these trickier classes, we can better elucidate differences between models. To this end, we further report EMNIST test error on only the ten hardest classes in Table 6. These correspond to handwritten digits and letters that are difficult to distinguish without context [^6] Specifically, the classes “0” and “O”; “1”, “I”, and “L”; “9”, “g”, and “q”; and “F” and “f”.. Here, the advantage of GLL is more pronounced in the 2% training case.
Overall, the results show that using a GLL classification head consistently performs on par with or better than MLP or WNLL-based classifiers. The benefit of a GLL is particularly evident at low label rates. To complement these results, Section 4.6 shows that a GLL offers substantial robustness gains under adversarial perturbations.
4.5 Training and Inference Times
Batch Size Metric GLL MLP
[…]
Testing time / batch (s) $0308{\pm 0.0085} 0.0$$0024{\pm 0.0027}$
Testing error (%) 0.76 2.11
256 Training time / batch (s) $0685{\pm 0.0146} 0.0$$0061{\pm 0.0005}$
Testing time / batch (s) $0457{\pm 0.0108} 0.0$$0039{\pm 0.0005}$
Testing error (%) 1.33 1.69
512 Training time / batch (s) $1244{\pm 0.0280} 0.0$$0094{\pm 0.0007}$
Testing time / batch (s) $0881{\pm 0.0206} 0.0$$0072{\pm 0.0006}$
Testing error (%) 1.27 1.53
1024 Training time / batch (s) $2447{\pm 0.0606} 0.0$$0167{\pm 0.0012}$
Testing time / batch (s) $1556{\pm 0.0310} 0.0$$0139{\pm 0.0009}$
Testing error (%) 1.26 4.65
Table 7: Comparison of GLL and MLP performance across different batch sizes on MNIST. The architecture backbone is a four-layer CNN with two linear layers. We use identical training settings for both models. For time metrics, each entry is reported as mean $\pm$ standard deviation over five epochs of training. Testing error is reported in percentage points. The experiments were run on a T4 GPU.
Batch Size Metric GLL MLP
128 Training time / batch (s) $1537{\pm 0.0121} 0.0$$0597{\pm 0.0083}$
Testing time / batch (s) $0659{\pm 0.0068} 0.0$$0220{\pm 0.0026}$
Testing error (%) 0.84 0.86
256 Training time / batch (s) $0.2295{\pm 0.0193}$ $0.1157{\pm 0.0168}$
Testing time / batch (s) $1045{\pm 0.0162} 0.0$$0426{\pm 0.0060}$
Testing error (%) 0.94 1.04
512 Training time / batch (s) $0.3851{\pm 0.0362}$ $0.2292{\pm 0.0323}$
Testing time / batch (s) $1834{\pm 0.0236} 0.0$$0924{\pm 0.0102}$
Testing error (%) 0.97 1.34
1024 Training time / batch (s) $0.7129{\pm 0.0773}$ $0.4676{\pm 0.0734}$
Testing time / batch (s) $0.3404{\pm 0.0352}$ $0.2032{\pm 0.0340}$
Testing error (%) 0.94 2.46
Table 8: Comparison of GLL and MLP performance across different batch sizes on MNIST. The architecture backbone is ResNet-18. We use identical training settings for both models. For time metrics, each entry is reported as mean $\pm$ standard deviation over five epochs of training. Testing error is reported in percentage points. The experiments were run on a T4 GPU.
Batch Size Metric GLL MLP
128 Training time / batch (s) $1257{\pm 0.0044} 0.0$$0471{\pm 0.0024}$
Testing time / batch (s) $0629{\pm 0.0035} 0.0$$0224{\pm 0.0037}$
Testing error (%) 1.65 2.39
256 Training time / batch (s) $1772{\pm 0.0065} 0.0$$0725{\pm 0.0028}$
Testing time / batch (s) $0950{\pm 0.0103} 0.0$$0404{\pm 0.0043}$
Testing error (%) 2.10 1.18
512 Training time / batch (s) $0.2950{\pm 0.0218}$ $0.1480{\pm 0.0637}$
Testing time / batch (s) $1605{\pm 0.0143} 0.0$$0774{\pm 0.0074}$
Testing error (%) 1.14 1.13
1024 Training time / batch (s) $0.5214{\pm 0.0605}$ $0.2802{\pm 0.0224}$
Testing time / batch (s) $0.2919{\pm 0.0208}$ $0.1498{\pm 0.0102}$
Testing error (%) 1.93 1.57
Table 9: Comparison of GLL and MLP performance across different batch sizes on MNIST. The architecture backbone is ResNet-110. We use identical training settings for both models. For time metrics, each entry is reported as mean $\pm$ standard deviation over five epochs of training. Testing error is reported in percentage points. The experiments were run on an A100 GPU.
To emphasize the practicality of our method, we present a side-by-side comparison of training and inference times of several GLL- and MLP-based models on the MNIST dataset in Tables 7, 8, and 9. We train each model for 5 epochs with varying batch sizes and report the average time per batch in training and testing. For GLL, we use a base labeled set of 100 points (10 per class). All other training hyperparameters are identical throughout. To ensure a fair comparison, we also report test error alongside training and inference time. Across all architectures and batch sizes, GLL attains test errors that are comparable to or better than those of the corresponding MLP baselines. For the small CNN backbone, GLL consistently achieves noticeably lower error than MLP, particularly at large batch sizes. For ResNet-18 and ResNet-110, the two methods are typically within a fraction of a percentage point of each other, with GLL often slightly better and only occasionally slightly worse. Note that this setting is simplistic - with no pretraining or hyperparameter optimization - and only meant to illustrate a computational efficiency comparison between two methods with similar test performance.
GLL is naturally slower due to the additional linear solve in the forward and backward passes: depending on the architecture and batch size, the per-batch training time increases by roughly a factor of two to an order of magnitude relative to MLP, with similar trends for testing time. Importantly, however, the extra cost of GLL is independent of the number of parameters in the backbone; its main overhead scales with the size of the base labeled set rather than network width or depth. Consequently, as the architecture becomes larger, from a small CNN to ResNet-18 and ResNet-110, GLL-based models progressively approach the runtime of their standard MLP-based counterparts, while retaining competitive or improved test error.
@@ 4.6 Adversarial Robustness
In this subsection, we demonstrate our method’s superior robustness to adversarial attacks compared to the typical projection head and softmax activation layer in a neural network. Adversarial attacks seek to generate adversarial examples (szegedy2013intriguing): input data with small perturbations that cause the model to misclassify the data. Many highly performant deep neural networks can be fooled by very small changes to the input, even when the change is imperceptible to a human (madry2017towards; goodfellow2014explaining). Hence, training models robust (i.e. resistant) to such attacks is important for both security and performance. Our experiments show that in both naturally and robustly-trained models, using the graph learning layer improves adversarial robustness compared to a softmax classifier on MNIST (mnist), FashionMNIST (xiao2017fashion), and CIFAR-10 (krizhevsky2009learning) in both relatively shallow and deep networks. Moreover, the GLL is more robust to adversarial attacks without sacrificing performance on natural images.
@@ 4.6.1 Adversarial Attacks
We consider three types of adversarial attacks: the fast gradient sign method (FGSM) (goodfellow2014explaining) in the $\ell_{\infty}$ norm, the iterative fast gradient sign method (IFGSM) (kurakin2018adversarial) in the $\ell_{\infty}$ norm, and the Carlini-Wagner attack (carlini2017towards) in the $\ell_{2}$ norm.
@@
where $x_{0}=x$ and i=0,,N1i=0,...,N-1, and so $x_{N}$ is the generated adversarial example. However, we must also clip the iterates (pixel-wise) during this iteration to ensure that (1) they remain in the $\epsilon$-neighborhood of $x$ and (2) they remain in the range of the image space (e.g. in [0,1] for a grayscale image). Thus, the IFGSM attack is:
@@
Note that this assumes the pixel values are in the range $[0,1]$, but can be easily tailored to other settings. IFGSM attacks have been shown to be more effective than FGSM attacks (kurakin2018adversarial), however we will see in Section 4.6.3 that our graph learning layer is highly robust to this iterative method.
@@
The constraint $C(x+\delta)=t$ is highly non-linear, so Carlini and Wagner instead consider objective functions $g$ that satisfy $C(x+\delta)=t$ if and only if $g(x+\delta)\leq 0$. While there are many choices for $g$ (see carlini2017towards, Section IV A), we will use the following in our experiments:
@@
where $a^{+}=\max(0,a)$. It is worth noting that a common choice for $g$ in the literature is to use the unnormalized logits from the neural network (that is, the output of a neural network just before the standard softmax layer) instead of $F(x)$. While both definitions satisfy the condition, we use the above because graph learning methods (such as Laplace learning) operate in the feature space and directly output probability distributions over the classes, so there is no notion of logits.
@@
where $c>0$ is a suitably chosen hyperparameter. To make the problem unconstrained, we can introduce a variable $w$ and write $\delta=\frac{1}{2}(\tanh{w}+1)-x$; since $1\leq\tanh{w}\leq 1$, we have that $0\leq x+\delta\leq 1$. This gives the unconstrained objective:
@@
which can be solved using standard optimization routines. We use the Adam optimizer (kingma2014adam).
@@ 4.6.2 Adversarial Training
To train adversarially robust networks, we use projected gradient descent (PGD) adversarial training (madry2017towards) on our models. Motivated by guaranteeing robustness to first-order adversaries, PGD training applies IFGSM to images with a random initial perturbation during training. Hence, the model is trained on adversarial examples. More precisely, PGD training consists of two steps each batch: first, $x$ is randomly perturbed
@@
where $x_{0}=x^{*}$ and the number of iterations and $\alpha$ are hyperparameters. This training procedure, while more expensive, has been shown to be effective at improving robustness in deep learning models compared to training on clean (i.e. natural) images (madry2017towards, wang2021graph).
@@ 4.6.3 Adversarial Experiments
We evaluate the adversarial robustness of neural networks with the GLL classifier on the MNIST (mnist), FashionMNIST (xiao2017fashion), and CIFAR-10 (krizhevsky2009learning) benchmark datasets, and compare our method to the same networks trained with either a softmax classifier or the WNLL classifier (wang2021graph). For all datasets, we normalize the images to have mean zero and standard deviation one, and we train and run all adversarial attacks on a 11GB GeForce RTX 2080Ti GPU. We note that the goal of these experiments is not to get state-of-the-art natural accuracies, but rather to demonstrate our method’s superior robustness to adversarial attacks compared to the typical MLP classifier without sacrificing performance on natural images. A summary of our experiments, highlighting a handful of our adversarial robustness results, can be found in Table 10, while complete results are shown in Figures 6, 7, and 8. Throughout, “MLP” refers to a model with a final linear layer mapping from the feature space to the logits followed by a softmax classifier, “WNLL” refers to a model with the WNLL classifier, and “GLL” refers to a model with our graph learning layer classifier. WNLL and GLL are both drop-in replacements for the MLP classifier, and applied on the feature space itself. For these experiments, we do not include any unlabeled samples $\mathcal{B}_{u}$ in batch training (because the setting is fully supervised), and we do not pretrain to demonstrate that GLL-based models can be trained from scratch with desireable natural and robust accuracies.
@@
MNIST GL Robust 0.55 2.28 2.30 1.16
WNLL Robust 0.58 2.48 2.97 2.72
MLP Robust 0.80 4.83 6.43 2.59
GL Natural 0.47 3.49 3.56 1.29
WNLL Natural 0.50 3.54 4.64 3.19
MLP Natural 0.75 5.74 8.48 2.76
Fashion-MNIST GL Robust 8.96 11.93 12.83 16.19
WNLL Robust 8.11 10.97 11.31 18.55
MLP Robust 7.92 13.50 14.33 19.00
GL Natural 7.86 34.56 38.19 42.81
WNLL Natural 6.44 21.46 39.33 74.59
MLP Natural 8.35 59.19 81.70 37.99
CIFAR-10 GL Robust 9.40 19.50 20.38 21.91
WNLL Robust 9.34 25.22 28.36 50.35
MLP Robust 8.71 26.05 28.96 24.06
GL Natural 5.71 40.95 45.59 22.77
WNLL Natural 5.13 58.20 92.33 93.87
MLP Natural 5.02 49.25 91.79 47.00
@@
Table 10: Summary of adversarial results of models trained with our method compared to WNLL and the standard MLP classifier, on both robustly and naturally trained models. We report natural error rate (i.e. no attacks), and adversarial error rate on FGSM, IFGSM, and CW attacks. Best error rates are in bold. For MNIST, we report $\epsilon=0.3$ for the FGSM and IFGSM attacks, and $\epsilon=0.05$ for the other datasets. We report CW results for $c=20$. We report more complete results across a range of values for $\epsilon$ and $c$ in Figures 6, 7, and 8. We include the natural error rates to show that the GLL does not sacrifice performance in this regime, while also being significantly more robust to adversaries across datasets. One mild exception is WNLL slightly outperforms GLL for the fast gradient attacks on FashionMNIST, however Figure 7 shows that in the case of IFGSM, GLL is much better against stronger attacks. The slight increase in natural test error on robustly trained models compared to naturally trained models has been observed previously in madry2017towards; however, as expected, the robust models perform significantly better on adversarially-perturbed data than their naturally-trained counterparts.
@@
For MNIST, we use the same “small CNN” architecture as in wang2021graph: four convolutional layers with ReLU activations after each layer and maxpooling after the second and fourth layer, followed by three linear layers and a softmax classifier. We train for 100 epochs using the Adam optimizer (kingma2014adam) with an initial learning rate of $0.01$ which decays by a factor of $0.1$ every 25 epochs. For the robustly trained models, we run PGD training with $\epsilon=0.3$ and $\alpha=0.01$ for 5 iterations. In training, we use a batch size of 1000, and for GLL and WNLL we use 100 labeled “base” points per batch (i.e. 10 per class), which are randomly sampled each epoch.
@@
To evaluate experimental robustness, we attack the test set with the FGSM, IFGSM, and CW attacks. When using FGSM and IFGSM on GLL, we embed the entire test set of 10000 digits along with 10000 randomly selected data points from the training set (i.e. 1000 per class) when running the attacks. We run CW in batches of 1000 test and 1000 training points due to computational constraints; in general graph learning methods perform better with more nodes on the graph as it better approximates the underlying data manifold. We use the same testing batch sizes (full batch on FGSM and IFGSM, 1000 for CW) for the MLP and WNLL models, and the same number of base points for WNLL. We report results for FGSM and IFGSM at varying levels of $\epsilon$, fixing $\alpha=0.05$ for IFGSM and letting the number of iterations be $5*(\epsilon/\alpha)$, following a similar heuristic as in kurakin2018adversarial to allow the adversarial example sufficient ability to reach the boundary of the $\epsilon$-ball. For CW attacks, we let $c$ vary and run 100 iterations of the Adam optimizer with a learning rate of 0.005. Results for each attack, comparing both robustly and naturally trained models with MLP and GLL classifiers, are shown in Figure 6.
@@
Figure 6: Adversarial robustness to FGSM, IFGSM, and CW attacks on the MNIST dataset. The backbone network is the same in all cases; the only difference is the classifier (either MLP with softmax, WNLL, or our novel GLL). We report results for both robustly and naturally trained models.
For FashionMNIST, we use ResNet-18 (he2016deep), where again we replace the final linear layer and softmax with GLL. We again use the Adam optimizer for 100 epochs, with an intial learning rate of $0.01$ which decays by a factor of $0.5$ every 10 epochs. For the robustly trained models, we run PGD training with $\epsilon=0.05$ and $\alpha=0.01$ for 5 iterations. In training, we use a batch size of 2000, and for GLL and WNLL we use 200 labeled “base” points per batch. We attack in batches of 250 test points, with 500 training points for GLL and WNLL; this was chosen as it was the largest batch size we could use given our computational resources. The hyperparameters for the attacks are identical to those for MNIST. We report results in Figure 7.
@@
Figure 7: Adversarial robustness to FGSM, IFGSM, and CW attacks on the FashionMNIST dataset. The backbone network is the same in all cases (ResNet-18); the only difference is the classifier (either MLP with softmax, WNLL, or our novel GLL). We report results for both robustly and naturally trained models.
Finally, for CIFAR-10, we use PreActResNet-18 (he2016identity) as our architecture backbone. We use stochastic gradient descent with momentum set to $0.9$, a weight decay of $0.0005$, and an intial learning rate of $0.1$. We train for 150 epochs and use cosine annealing to adjust the learning rate (loshchilov2016sgdr). For the robustly trained models, we run PGD training with $\epsilon=0.05$ and $\alpha=0.01$ for 5 iterations. In training, we use a batch size of 200, and for GLL and WNLL we use 100 labeled “base” points per batch. We attack in batches of 200 test points, with 500 training points for GLL and WNLL; this was chosen as it was the largest batch size we could use given our computational resources. The hyperparameters for the attacks are identical to those for MNIST, except we use 50 iterations of optimization for the CW attacks. We report results in Figure 8.